STEP-SCAN FOURIER TRANSFORM INFRARED ABSORPTION SPECTROSCOPY OF CUBIC SOLID ACETYLENE CLUSTERS

\underline{SHINICHI HIRABAYASHI} and YASUHIRO HIRAHARA, Division of Earth and Environmental Sciences, Graduate School of Environmental Studies, Nagoya University, Furo-cho, Chikusa-ku, Nagoya 464-8602, Japan.

The infrared spectra of large acetylene clusters formed in pulsed supersonic jets have been measured in the 800-5000 cm^{−1}region using step-scan Fourier transform infrared (FTIR) absorption spectroscopy.^{a} The ν_{3} band of C_{2}H_{2} cluster was reproduced well by two Lorentzian profiles peaked at 3234 and 3241 cm^{−1}. From comparison with the spectral positions of small clusters and two solid phases in the C-H stretching region, the 3234 cm^{−1}band was assigned to the solid in the high temperature cubic phase. Another feature is invariant at 3241 cm^{−1}under various supersonic jet conditions and is therefore likely attributed to the large clusters with stable structure. The relative intensity for monomer vibrationrotation lines gives the rotational temperature of 12 K at 10% C_{2}H_{2} in Ar. The infrared absorption band of C_{2}D_{2} clusters has also been measured at 2399 cm^{−1}. The deuterium isotopic shift is in good agreement with the predictions based on the harmonic oscillator model and confirms the assignment of cubic C_{2}H_{2} and C_{2}D_{2}. This is the first infrared observation of cubic solid C_{2}D_{2}. Although the experimental data do not allow us to estimate the precise size of cubic acetylene clusters, the cluster size is likely to be ∼10 nm.

aS. Hirabayashi and Y. Hirahara, Chem. Phys. Lett. 361, 265 (2002).